# IslamicMMLU: A Benchmark for Evaluating LLMs on Islamic Knowledge

### 24 Mar 2026 v1

@@ A Benchmark for Evaluating LLMs on Islamic Knowledge
Ali Abdelaal, Mohammed Nader Al Haffar,
Mahmoud Fawzi, Walid Magdy
The University of Edinburgh {s2431177, s2419871, m.fawzi}@ed.ac.uk, wmagdy@inf.ed.ac.uk
@@ 1 Introduction
Large language models (LLMs) are increasingly consulted for information about Islamic topics by millions of users worldwide Mubarak et al. (2025). Islamic knowledge spans multiple interconnected disciplines (e.g., Quranic studies, Hadith sciences, and Islamic jurisprudence), each requiring distinct scholarly expertise and evaluation methodologies. Yet no comprehensive benchmark evaluates LLM performance across these core domains. This gap can lead to religious misinformation affecting daily practice for Muslims who may lack easy access to qualified scholars (Naous et al., 2024; Fawzi et al., 2026).
@@ 2.1 Islamic Knowledge Domains
The Hadith literature records the sayings, actions, and tacit approvals of the Prophet Muhammad (Fawzi et al., 2025). It consists of two parts: 1) isnad which is the chain of narrators who narrated the quote of the prophet till the time it has been written, and 2) matn, which is the main content of the quote itself. Islam has two main branches. The Sunni branch which is followed by the majority of Muslims and the Shia branch. Six canonical collections of Hadith are widely accepted by Sunni Muslims as the most reliable source of Hadith, which are Bukhari, Muslim, Abu Dawud, Tirmidhi, Nasa’i, and Ibn Majah. Hadith sciences classify narrations by authenticity according to how evident it has been said by the prophet. There are four main authenticity levels: sahih (authentic), hasan (good), da‘if (weak), and mawdu‘ (fabricated). Shia Muslims use other books.
@@ 2.3 Arabic NLP Benchmarks
ArabicMMLU (Koto et al., 2024) is the most direct comparator: 14,575 native Arabic questions across 40 subjects sourced from school exams in eight Arab countries, with quality validation on 100 samples achieving 96% accuracy. It includes Islamic studies as a single category with basic questions without the domain depth needed to examine knowledge in different nuances of Islam. ACVA, introduced with AceGPT (Huang et al., 2024), offers 8,000+ true/false questions on Arabic cultural values but explicitly excludes religious jurisprudence. AlGhafa (Almazrouei et al., 2023) provides a general Arabic evaluation benchmark but does not cover Islamic knowledge domains in depth. ILMAAM (Nacar et al., 2025) provides a culturally aligned Arabic MMLU variant. ArabLegalEval (Hijazi et al., 2024) evaluates LLMs on Saudi secular law with 10,000+ MCQs for domain-specific evaluation. A recent survey catalogues over 40 Arabic benchmarks and identifies gaps in religious domain coverage (Al-Zubaidi et al., 2025).
@@ 2.4 Cultural Bias in LLMs
Naous et al. (2024) demonstrate that LLMs default to Western cultural norms even when prompted in Arabic. Plaza-del Arco et al. (2024) find differential treatment across faiths, with Islam facing elevated stereotyping and disproportionate refusal rates. Abid et al. (2021) provide foundational evidence of persistent anti-Muslim bias in GPT-3. DLAMA (Keleg and Magdy, 2023) reveals Western bias in multilingual knowledge probing.
@@ 2.5 Islamic and Religious NLP
PalmX (Al-Wajih et al., 2025) addresses cultural competence including an Islamic culture subtask. QIAS (Bouchekif et al., 2025) covers Islamic inheritance reasoning. FiqhQA (Atif et al., 2025) introduces 960 question-answer pairs categorised by school but relies on synthetically generated answers from a single LLM, raising concerns about circular evaluation. IslamicLegalBench (Elmahjub et al., 2026) provides 718 manually curated instances across seven schools of jurisprudence, measuring legal reasoning depth. QuranQA (Malhas and Elsayed, 2022) establishes reading comprehension benchmarks for Quranic text. IslamTrust (Lahmar et al., 2025) evaluates alignment with consensus Islamic ethics, achieving only 66.5% alignment with GPT-4.

### 03 Apr 2026 v2

@@ A Benchmark for Evaluating LLMs on Islamic Knowledge
Ali Abdelaal, Mohammed Nader Al Haffar, Mahmoud Fawzi, Walid Magdy The University of Edinburgh {A.Abdelaal, M.N.Al-haffar, M.F.G.Ibrahim}@sms.ed.ac.uk wmagdy@inf.ed.ac.uk
@@ 1 Introduction
Large language models (LLMs) are increasingly consulted for information about Islamic topics by millions of users worldwide Mubarak et al. (2025). Islamic knowledge spans multiple interconnected disciplines (e.g., Quranic studies, Hadith sciences, and Islamic jurisprudence), each requiring distinct scholarly expertise and evaluation methodologies. Yet no comprehensive benchmark evaluates LLM performance across these core domains. This gap can lead to religious misinformation affecting daily practice for Muslims who may lack easy access to qualified scholars (Naous et al., 2024; Fawzi et al., 2025b).
@@ 2.1 Islamic Knowledge Domains
The Hadith literature records the sayings, actions, and tacit approvals of the Prophet Muhammad (Fawzi et al., 2025a). It consists of two parts: 1) isnad which is the chain of narrators who narrated the quote of the prophet till the time it has been written, and 2) matn, which is the main content of the quote itself. Islam has two main branches. The Sunni branch which is followed by the majority of Muslims and the Shia branch. Six canonical collections of Hadith are widely accepted by Sunni Muslims as the most reliable source of Hadith, which are Bukhari, Muslim, Abu Dawud, Tirmidhi, Nasa’i, and Ibn Majah. Hadith sciences classify narrations by authenticity according to how evident it has been said by the prophet. There are four main authenticity levels: sahih (authentic), hasan (good), da‘if (weak), and mawdu‘ (fabricated). Shia Muslims use other books.
@@ 2.3 Arabic NLP Benchmarks
ArabicMMLU (Koto et al., 2024) is the most direct comparator: 14,575 native Arabic questions across 40 subjects sourced from school exams in eight Arab countries, with quality validation on 100 samples achieving 96% accuracy. It includes Islamic studies as a single category with basic questions without the domain depth needed to examine knowledge in different nuances of Islam. ACVA, introduced with AceGPT (Huang et al., 2024), offers 8,000+ true/false questions on Arabic cultural values but explicitly excludes religious jurisprudence. AlGhafa (Almazrouei et al., 2023) provides a general Arabic evaluation benchmark but does not cover Islamic knowledge domains in depth. ILMAAM (Nacar et al., 2025) provides a culturally aligned Arabic MMLU variant. ArabLegalEval (Hijazi et al., 2024) evaluates LLMs on Saudi secular law with 10,000+ MCQs for domain-specific evaluation. A recent survey catalogues over 40 Arabic benchmarks and identifies gaps in religious domain coverage (Alzubaidi et al., 2025).
@@ 2.4 Cultural Bias in LLMs
Naous et al. (2024) demonstrate that LLMs default to Western cultural norms even when prompted in Arabic. Plaza-del-Arco et al. (2024) find differential treatment across faiths, with Islam facing elevated stereotyping and disproportionate refusal rates. Abid et al. (2021) provide foundational evidence of persistent anti-Muslim bias in GPT-3. DLAMA (Keleg and Magdy, 2023) reveals Western bias in multilingual knowledge probing.
@@ 2.5 Islamic and Religious NLP
PalmX (Alwajih et al., 2025) addresses cultural competence including an Islamic culture subtask. QIAS (Bouchekif et al., 2025) covers Islamic inheritance reasoning. FiqhQA (Atif et al., 2025) introduces 960 question-answer pairs categorised by school but relies on synthetically generated answers from a single LLM, raising concerns about circular evaluation. IslamicLegalBench (Elmahjub et al., 2026) provides 718 manually curated instances across seven schools of jurisprudence, measuring legal reasoning depth. QuranQA (Malhas and Elsayed, 2022) establishes reading comprehension benchmarks for Quranic text. IslamTrust (Lahmar and others, 2025) evaluates alignment with consensus Islamic ethics, achieving only 66.5% alignment with GPT-4.